OpenStudy (anonymous):

ΔABC and ΔGHI are similar triangles. What is the perimeter of ΔGHI? can someone help will give medal

OpenStudy (anonymous):

the image woont show

OpenStudy (anonymous):

here it is

OpenStudy (anonymous):

80/20 = 96/x cross multiply,, x= (GI) =24 perimeter = 24+20+14 = 58 ?

OpenStudy (anonymous):

i get it thanks @yajna

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

u must use similar triangles